# Neural ODEs for holographic transport models without translation symmetry

### v1

@@ Abstract
We investigate the data-driven holographic transport models without translation symmetry. Our data are chosen as the real part of the frequency-dependent shear viscosity. We develop a radial flow equation for a large class of holographic models, which determine the shear viscosity by the black hole metric and the graviton mass. The latter serves as the bulk dual of the translation symmetry breaking on the boundary. We convert the flow equation to a Neural Ordinary Differential Equation (Neural ODE), which is a neural network with continuous depth and produces output through a black-box ODE solver. Given either the metric or the mass, we illustrate that the Neural ODE can learn the other with high accuracy. Our work demonstrates the capabilities of Neural ODEs in bulk reconstruction and applied holography.
@@ 1 Introduction
AdS/CFT correspondence is an elegant holographic mapping that brings illuminating insights into quantum gravity and provides fruitful phenomenal models of strongly coupled quantum field theories [1, 2]. The typical process for the holographic modeling involves assuming a bulk action, solving the Einstein equation to obtain the bulk spacetime metric, and using the holographic dictionary to convert various gravity observables into field-theory observables. However, building a suitable action from a phenomenological perspective is a fundamental and challenging inverse problem in theoretical physics. In addition to relying on the symmetry principle, it usually requires deep domain knowledge and intuition from human experts.
@@
In holography, under certain conditions, some field-theory observables can be determined solely by the bulk metric, without requiring the full action. This leads to a reduced inverse problem: Can we reconstruct the bulk metric from these field-theory observables? This line of work belongs to an important branch of the well-known bulk reconstruction project [4, 3, 5]. Its ambitious goal is to reorganize the CFT degrees of freedom into local gravitational physics under certain limits. There are many different methods for reconstructing the bulk metric, part of which can be found in [6, 7, 8, 9, 12, 10, 11, 13, 14, 15, 16, 17, 18, 19, 20, 22, 21, 23, 24, 25, 26]. Among others, Hashimoto et al. propose an interesting scheme using the neural network and deep learning (DL) [27]. They discretize the Klein-Gordon equation of a scalar field in a curved spacetime as the architecture of a neural network. Their data consist of the VEV of the operator and its conjugate source, with the IR behavior of the field as the labels. They use a smoothness term on the metric as a regularization to guide the training. After training, the black hole metric is encoded as the weights of the neural network. This scheme is called as the AdS/DL correspondence. It has been quickly applied to construct holographic QCD models and obtained interesting results, which include extracting the novel feature from the lattice data of chiral condensate, predicting the reasonable mass for excited state mesons, and deriving the explicit form of a dilaton potential, etc [28, 29, 30, 31, 32, 33].
@@
In addition to QCD, holographic models have been widely used in condensed matter physics, hydrodynamics, and quantum information. Among others, transport phenomena, such as momentum and charge transports, are considered very suitable for the holographic research based on the bottom-up setup. In view of this, Yan et al. propose a data-driven holographic transport model by deep learning [34]. Their data are the complex frequency-dependent shear viscosity and the network architecture corresponds to the discretization of a radial flow equation. Since the IR solution of the flow equation can be fully determined by regularity, the forward propagation of information is designed from IR to UV. Compared with previous algorithms, this algorithm eliminates the need to introduce additional IR labels for the data, thereby reducing systematic errors in the analysis. Recently, a similar algorithm has been developed to learn from the optical conductivity [35], aiming to apply AdS/DL to the condensed matters, such as strange metal and high-temperature superconductivity.
@@
All of the above work about AdS/DL has one commonality: the emergent extra dimension corresponds to the depth of the neural network. In fact, the neural network itself is regarded as a bulk spacetime. However, since the neural network has discrete layers, both the spacetime metric and the equation of motion are discrete. This leads to intrinsic discrete errors in the algorithm. Moreover, the regularization term that induces smooth metrics increases arbitrariness and computational cost.
@@
To address these discrete problems, Neural ODEs have been introduced for AdS/DL [36]. This is a relatively new family of deep neural network models that generalize the standard layer-to-layer propagation to continuous-depth models [37]. In essence, Neural ODEs employ a neural network to parameterize the derivative of the hidden state, while the output is computed using a black-box adaptive differential equation solver. The application of Neural ODEs in holographic modeling, as highlighted in [36], showcases the potential of AI as more than just a data processing tool in theoretical physics, but also as a valuable aid in scientific discoveries.
The purpose of this paper is twofold. On a technical level, we aim to replace the discrete architecture in [34] with Neural ODEs. On a physical level, we intend to break the translation symmetry that is respected by the theoretical framework in [34]. The latter is indispensable to describe the real-world condensed matter. However, for AdS/DL to be significant, its theoretical framework for the holographic calculation must possess a certain universality. In [34], the universality depends heavily on the relationship between the shear viscosity of the boundary field theory and the shear response on the UV boundary. If the translation symmetry is broken, this relationship usually no longer holds. Fortunately, through the analysis of the UV behavior related to the holographic renormalization [38, 39], we find that the shear viscosity and the shear response on the boundary usually have the same real part, even if the translation symmetry is broken. This observation motivates us to use only the real part of the shear viscosity as the data for learning.
After the data is ready, we need to clarify the training tasks. In this paper, we will focus on the holographic models with homogeneous and isotropic geometry, which are relatively simple. Nevertheless, determining the shear viscosity requires the function of graviton mass in addition to the function of metric. As a reverse engineering, we will study if given a mass or a metric, can Neural ODEs learn the other one? [^1] Since there is a trade-off between metric and mass in determining shear viscosity, we do not attempt to learn the two functions simultaneously.
@@
The rest of the paper will be arranged as follows. In Section 2, we will develop a radial flow equation in a large class of holographic models without translation symmetry. In Section 3, we will introduce the machine learning algorithm, including the network architecture, loss function and inductive bias. In Section 4, we will generate the data and depict the training results based on three typical holographic models without translation symmetry. The conclusion and discussion will be presented in Section 5. There are also two appendices. In Appendix A, we will build the relationship between the shear viscosity and the shear response. In Appendix B, we will provide the training scheme and report.
@@ 2 Holographic flow equation
Importantly, Eq. (1) and Eq. (2) are homogeneous and isotropic along the field theory directions, although we have not assumed that the matter fields are homogeneous. When the black hole is perturbed by the time-dependent shear mode $\left(\delta g\right)_{\;x_{2}}^{x_{1}}=h(r)e^{-i\omega t}$, the wave equation has a general form [40]
@@
where $m$ is the radially varying graviton mass, given by
$m\left(r\right)^{2}=g^{xx}T_{xx}-\frac{\delta T_{xy}}{\delta g_{xy}}.$ (4)
Note that the nonzero graviton mass in the bulk is dual to the breaking of the translation symmetry on the boundary [41].
@@
where $r_{\mathrm{h}}$ is the horizon radius. Using Eq. (7) as the IR boundary condition, the flow equation can be solved to obtain the complex shear response at UV. In general, however, it is not the desired shear viscosity. Nevertheless, by the UV analysis given in Appendix A, we will demonstrate that for a large class of theories, their real parts are the same indeed
$\eta_{\mathrm{re}}=\left.\chi_{\mathrm{re}}\right|_{r\rightarrow\infty}.$ (8)
Now, let’s simplify the metric (1) by imposing $g_{tt}g_{rr}=1$ and setting $g_{xx}=r^{2}$. Then one can obtain
$ds^{2}=\frac{1}{z^{2}}\left[-f(z)dt^{2}+\frac{1}{f(z)}dz^{2}+d\vec{x}^{2}\right],$ (9)
where the coordinate $z=r_{\mathrm{h}}/r$ has been used. Thus, the horizon is located at $z_{\mathrm{h}}=1$ and the boundary at $z_{\mathrm{b}}=0$. In terms of Eq. (9), the flow equation (6) can be reduced to
$\partial_{z}\chi+\frac{i\omega}{f}\left(z^{2}\chi^{2}-\frac{1}{z^{2}}\right)-\frac{1}{i\omega}\frac{m^{2}}{z^{4}}=0,$ (10)
where the horizon radius has been set as $r_{\mathrm{h}}=1$. Accordingly, the IR boundary condition (7) and the UV relation (8) are changed into $\chi(z_{\mathrm{h}})=1$ and $\eta_{\mathrm{re}}=\chi_{\mathrm{re}}\left(z_{\mathrm{b}}\right)$, respectively.
@@ 3.1 Neural ODE
In residual networks [42], there is a sequence of transformations to a hidden state
@@
Given an initial state $x(t_{0})$, the forward propagation of Neural ODEs yields the state $x(t_{1})$ through a black-box ODE solver. The back propagation of Neural ODEs is realized by the adjoint sensitivity method [43]. Suppose that there is a loss function depending on the outputs of an ODE solver
@@
and the parameter gradient can be calculated by an integral [37]
@@
Then we identify the shear response $\left(\chi_{\mathrm{re}},\chi_{\mathrm{im}}\right)$ with the state $x$ and the radial coordinate $z$ with the time $t$. In general, the function of metric $f(z)$ or the function of mass square $m^{2}(z)$ can be represented by a neural network with the parameter $\theta$ [^2] In [36], the metric ansatz is a polynomial, which constrains the representation capability of Neural ODEs. In this work, we will eliminate this constraint..
Obviously, given the parameter $\theta$ is equivalent to giving the trial function of metric or mass square which specifies the Neural ODE. By invoking an ODE solver, one can integrate from IR to UV and predict the real part of shear viscosity. Using the adjoint sensitivity method, one can implement back-propagation and perform training. We depict the architecture of the current algorithm in Fig. 1. Note that it differs from the usual Neural ODE in that one component of the adjoint state vanishes.
@@ 3.2 Loss function and inductive bias
Our loss function is chosen as the log-cosh form [44, 45, 46, 47]
@@
where $N$ denotes the size of data, $\bar{\eta}_{\mathrm{re}}$ represents the input data, and $\eta_{\mathrm{re}}$ is what the Neural ODE predicts. Note that the loss $\log\left[\cosh\left(x\right)\right]$ behaves like L1 loss $\left|x\right|$ far from the origin and L2 loss $x^{2}$ close to the origin. It is twice differentiable everywhere, making it suitable for optimizers like BFGS that require the Hessian [48].
@@ 4 Holographic models
We will study three holographic models without translation symmetry, all of which have homogeneous and isotropic geometry.
@@ 4.1 Massive gravity
The research of massive gravity (MG) has a long and winding history [49, 50, 51, 52, 53]. The main theoretical interest is to explore whether there exists a self-consistent theory that has a massive graviton with spin 2. In cosmology, MG is considered as a candidate theory that explains the accelerated expansion of the universe by modifying Einstein gravity. In holography, massive gravity is the first analytically tractable model without translation symmetry [54]
We will consider the dRGT massive gravity that is claimed to be ghost free [55, 56, 57]. Its simplest version used in holography is described by the bulk action
@@
where $\mathcal{X}{{}^{\mu}}_{\nu}=\sqrt{g^{\mu\lambda}f_{\lambda\nu}}$ and $f_{\mu\nu}=\mathrm{diag}(0,0,1,1)$ is the reference metric. Note that we have set the gravitational constant $116\pi G=$ 1 and the AdS radius $L=1$. The Gibbons-Hawking term is given by
$S_{\mathrm{GH}}=2\int d^{3}x\sqrt{-\gamma}K,$ (23)
where $\gamma_{ij}$ is the induced metric and $K$ is the external curvature. The counterterms have been derived in [58] [^3] Although a clear form of counterterms is not necessary for this section, it can help the reader double check the analysis given in Appendix A.
$S_{\mathrm{ct}}=\int d^{3}x\sqrt{-\gamma}\left[-4-R+\frac{1}{2}\alpha e_{1}+\frac{1}{16}\alpha^{2}e_{2}+\frac{\alpha}{4}(2R_{ij}Y^{ij}-e_{1}R)\log r\right],$ (24)
where $e_{1}=\mathrm{tr}Y$, $e_{2}=\left(\mathrm{tr}Y\right)^{2}-\mathrm{tr}Y^{2}$, and $Y{{}^{i}}_{j}=\sqrt{\gamma^{ik}f_{kj}}$.
@@
Now we can generate the data for learning. Inputting Eq. (26) and Eq. (29) into Eq. (10), using the regular condition at the horizon, and setting the typical parameter $\alpha=1$, one can solve the flow equation and obtain the real part of shear viscosity on the boundary. We sample the frequency uniformly between 0.01 and 6 to generate 600 data points $\left[\omega,\eta_{\mathrm{re}}(\omega)\right]$, which are plotted in Fig. 2.
Figure 2: The real part of frequency-dependent shear viscosity. It is generated by solving the flow equation with the true metric and mass in the MG model.
We conduct two numerical experiments: fixing the mass to learn the metric and fixing the metric to learn the mass. The primary training results are shown in Fig. 3 below. Additionally, Fig. 8 and Tab. 1 in Appendix B detail the hyperparameter tuning, minimum loss, and mean relative error (MRE).
Figure 3: The performance of Neural ODEs for learning metric (left) and mass (right) in the MG model. The curves represent the true values and the dots indicate the predictions.

### v2

@@ Abstract
We investigate the data-driven holographic transport models without translation symmetry. Our data are the real part of the frequency-dependent shear viscosity, denoted as $\eta_{\mathrm{re}}(\omega)$. We develop a radial flow equation of the shear response and establish its relation to $\eta_{\mathrm{re}}(\omega)$ for a wide class of holographic models. This allows us to determine $\eta_{\mathrm{re}}(\omega)$ of a strongly coupled field theory by the black hole metric and the graviton mass. The latter serves as the bulk dual to the translation symmetry breaking on the boundary. We convert the flow equation to a Neural Ordinary Differential Equation (Neural ODE), which is a neural network with continuous depth and produces output through a black-box ODE solver. We test the Neural ODE in three well-known holographic models without translation symmetry. Given either the metric or the mass, we display that the Neural ODE can learn the other with high accuracy. We also illustrate that the learned metric can be used to predict the derivative of entanglement entropy $S$ with respect to the size of entangling region $l$, which suggests an intriguing relationship between $S^{\prime}(l)$ and $\eta_{\mathrm{re}}(\omega)$.
@@ 1 Introduction
The AdS/CFT correspondence is an elegant holographic mapping that brings illuminating insights into quantum gravity and provides fruitful phenomenal models of strongly coupled quantum many-body systems Aharony9905 ; Liu2004 . The standard approach to the holographic modeling involves positing a bulk action, solving the Einstein equation to derive the bulk metric, and then using the holographic dictionary to transform various gravitational observables into field-theory observables. However, constructing a suitable action from a phenomenological perspective is a fundamental and challenging inverse problem in theoretical physics. In addition to relying on the symmetry principle, it usually requires deep domain knowledge and intuition from human experts.
@@
In certain conditions within the realm of holography, some field-theory observables can be derived directly from the bulk metric without necessitating the complete action. This leads to a reduced inverse problem: Can we reconstruct the bulk metric from these field-theory observables? This question is central to an important branch of the broader bulk reconstruction project Harlow:2018fse ; DeJonckheere:2017qkk ; Kajuri2003 , which ambitiously aims to reorganize the degrees of freedom in the CFT into local gravitational physics under specific limits. There are many different methods for reconstructing the bulk metric, part of which can be found in deHaro:2000vlm ; Hammersley:2006cp ; Hubeny:2006yu ; Hammersley:2007ab ; Hubeny:2012ry ; Bilson:2008ab ; Bilson:2010ff ; Balasubramanian:2013lsa ; Myers:2014jia ; Czech:2014ppa ; Engelhardt:2016wgb ; Engelhardt:2016crc ; Roy:2018ehv ; Kabat:2018smf ; Hashimoto:2020mrx ; Caron-Huot2211 ; Hashimoto:2021umd ; Jokela:2020auu ; Jokela2304 ; XWB2023 ; Qi2306 ; Yang2023 . Among others, Hashimoto et al. propose an intriguing scheme using the neural network and deep learning (DL) Hashimoto:2018ftp . They discretize the Klein-Gordon equation for a scalar field in curved spacetime into the architecture of a neural network. The data they use comprises the expectation value of the operator and its conjugate source, with the IR behavior of the field serving as labels. Under the inductive bias that the metric should be smooth, they design a regularization term to guide the training process. After training, the bulk metric is encoded as the weights of the neural network. This scheme is termed the AdS/DL correspondence. It has been quickly applied to construct holographic QCD models and produced interesting outcomes, such as extracting novel features from lattice data on chiral condensate, predicting reasonable masses for excited-state mesons, and deriving the explicit form of a dilaton potential Hashimoto:2018bnb ; Akutagawa:2020yeo ; Hashimoto:2021ihd .
@@
In addition to QCD, holographic models have found extensive applications in condensed matter physics, hydrodynamics, and quantum information. Notably, transport phenomena, such as momentum and charge transports, are considered very suitable for the holographic research based on the bottom-up setup. In this context, Yan et al. propose a data-driven holographic transport model Wu2020 . Their data are the complex frequency-dependent shear viscosity and the network architecture corresponds to the discretization of a radial flow equation. Since the IR solution of the flow equation is fully determined by regularity, the forward propagation of information is designed from IR to UV. Compared to previous algorithms, this design eliminates the need for additional IR labels, reducing systematic errors in the analysis. Recently, a similar algorithm has been developed to learn from optical conductivity Li:2022zjc , which has the potential applications in condensed matter physics, such as strange metal and high-temperature superconductivity.
@@
All the aforementioned works on AdS/DL share a common feature: the emergent extra dimension corresponds to the depth of the neural network. In fact, the neural network itself is regarded as a bulk spacetime. However, since the neural network has discrete layers, both the spacetime metric and the equation of motion are discretized. This leads to intrinsic discretization errors in the algorithm. Additionally, the regularization term used to induce smooth metrics increases both arbitrariness and computational cost.
@@
To address these discretization problems, Neural ODEs have been introduced for AdS/DL Hashimoto:2020 ; Hashimoto:2022eij . This is a relatively new family of deep neural network models that generalize the standard layer-to-layer propagation to continuous-depth models Chen:2018 . In essence, Neural ODEs employ a neural network to parameterize the derivative of the hidden state, while the output is computed using a black-box adaptive differential equation solver. The application of Neural ODEs in holographic modeling, as highlighted in Hashimoto:2020 , showcases the potential of AI as more than just a data processing tool in theoretical physics, but also as a valuable aid in scientific discoveries.
The purpose of this paper is twofold. On a technical level, we aim to replace the discrete architecture in Wu2020 with Neural ODEs. Note that the metric ansatz was limited to a polynomial in Hashimoto:2020 ; Hashimoto:2022eij , which constrains the representation capability of Neural ODEs. In this work, we will eliminate this constraint by incorporating the neural network into the ansatz. It should be pointed out that switching from polynomials to neural networks increases the training difficulty significantly. Fortunately, we find an effective training method. The key strategies include using three optimizers (RMSProp, Adam, BFGS) in sequence, as well as employing a larger number of initial values and the early-stopping technique to prevent Neural ODEs from getting stuck with bad parameters. On a physical level, we intend to break the translation symmetry that is respected by the theoretical framework in Wu2020 . The latter is indispensable to describe the real-world condensed matter. However, the breaking of translation symmetry presents an important challenge. Following the spirit of AdS/DL, one would expect the framework of the theory to be as universal as possible, so that the machine has the potential to learn novel metrics from the data of field theories, where the gravity dual may not be explicit. In Hashimoto:2018ftp , the universality is achieved under the probe limit, where the bulk action generating the background metric can be entirely unknown. In Wu2020 , the probe limit is not invoked, but the universality depends heavily on the relationship between the shear viscosity of the boundary field theory and the shear response on the UV boundary. If the translation symmetry is broken, this relationship usually no longer holds. In this work, through a careful analysis of the UV behavior related to the holographic renormalization Skenderis0209 ; Papadimitriou2016 , we find that the shear viscosity and the shear response on the boundary usually have the same real part, even if the translation symmetry is broken. This crucial observation motivates us to use only the real part of the shear viscosity as the training data.
Once the data is prepared, we will define our learning tasks. Following Hartnoll1601 , we will focus on the holographic models that break translation symmetry but maintain homogeneity and isotropy in geometry. In many of these models, we will demonstrate that the real part of the frequency-dependent shear viscosity is determined exclusively by functions of the bulk metric and the graviton mass. As a reverse engineering, our objective is to investigate whether Neural ODEs can learn one of these functions when the other is given [^1] Since there is a trade-off between metric and mass in determining shear viscosity, we do not attempt to learn the two functions simultaneously..
@@
The rest of this paper is organized as follows. In Section 2, we will develop a radial flow equation for a large class of the holographic models without translation symmetry. In Section 3, we will introduce the machine learning algorithm, including the architecture, loss function and inductive bias. In Section 4, we will generate the data and display the training results based on three well-known holographic models without translation symmetry. The conclusion and discussion will be presented in Section 5. Additionally, there are four appendices. In Appendix A, we will establish the relationship between the shear viscosity and the shear response. In Appendix B, we will provide the training scheme and report. In Appendix C, we will improve the performance of Neural ODEs at low temperatures. In Appendix D, we will study the entanglement entropy (EE) on the boundary using the bulk metric learned by Neural ODEs.
@@ 2 Holographic flow equation
Importantly, eq. (1) and eq. (2) are homogeneous and isotropic along the field theory directions, although we have not assumed that the matter fields are homogeneous. When the black hole is perturbed by the time-dependent shear mode $\delta g_{xy}=g_{xx}h(r)e^{-i\omega t}$, the wave equation has a general form Hartnoll1601  [^2] Here we have assumed that this mode decouples from other perturbations. Since $\delta g_{xy}$ transforms as a parity-even tensor mode in $x-y$ plane, this assumption is quite general. However, one can find a counterexample in the Einstein-$SU(2)$ theory Son1311 , where the coupling appears from a particular background that locks the symmetries in spacetime and $SU(2)$ vector space.
@@
where $m$ is the radially varying graviton mass
$m\left(r\right)^{2}=g^{xx}T_{xx}-\frac{\delta T_{xy}}{\delta g_{xy}},$ (4)
[…]
@@
where $r_{\mathrm{h}}$ is the horizon radius. Using eq. (7) as the IR boundary condition, the flow equation can be solved to obtain the complex shear response at UV. However, this shear response does not generally represent the desired shear viscosity on the boundary field theory. Nevertheless, through a careful UV analysis in Appendix A, we will demonstrate that for a wide range of theories, their real parts are indeed the same:
$\eta_{\mathrm{re}}\left(\omega\right)=\left.\chi_{\mathrm{re}}\left(\omega\right)\right|_{r\rightarrow\infty}.$ (8)
We emphasize that this nontrivial observation is crucial to the theoretical framework of our data-driven model.
To proceed, we will exploit the scaling symmetry of this system
$r\rightarrow r\lambda,\ \left(t,x,y\right)\rightarrow\left(t,x,y\right)/\lambda,\ g_{rr}\rightarrow g_{rr}\lambda^{-2},\ g_{tt}\rightarrow g_{tt}\lambda^{2},\ g_{xx}\rightarrow g_{xx}\lambda^{2},$ (9)
which indicates that one can set $r_{\mathrm{h}}=1$ for convenience.
We will further simplify the metric (1) by imposing $g_{tt}g_{rr}=1$ and setting $g_{xx}=r^{2}$. Then the metric ansatz can be written as
$ds^{2}=-r^{2}f(r)dt^{2}+\frac{1}{r^{2}f(r)}dr^{2}+r^{2}d\vec{x}^{2}.$ (10)
Replacing the radial coordinate with $z=1/r$, one can see that the horizon is located at $z_{\mathrm{h}}=1$ and the boundary at $z_{\mathrm{b}}=0$. Accordingly, the flow equation (6) can be reduced to [^3] If $g_{tt}g_{rr}\neq 1$, this equation is still valid, but $f$ and $m^{2}$ will represent the joint factors $\frac{1}{r^{2}}\sqrt{\frac{g_{tt}(r)}{g_{rr}(r)}}$ and $\sqrt{g_{tt}(r)g_{rr}(r)}m(r)^{2}$, respectively.
$\partial_{z}\chi+\frac{i\omega}{f(z)}\left(z^{2}\chi^{2}-\frac{1}{z^{2}}\right)-\frac{1}{i\omega}\frac{m^{2}}{z^{4}}=0,$ (11)
and the IR boundary condition (7) and the UV relation (8) are changed into $\chi(z_{\mathrm{h}})=1$ and $\eta_{\mathrm{re}}=\chi_{\mathrm{re}}\left(z_{\mathrm{b}}\right)$, respectively.
@@ 3.1 Neural ODE
In residual networks He1512 , there is a sequence of transformations to a hidden state
@@
Given an initial state $x(t_{0})$, the forward propagation of Neural ODEs yields the state $x(t_{1})$ through a black-box ODE solver. The back propagation of Neural ODEs is realized by the adjoint sensitivity method Pontryagin:1962 . Suppose that there is a loss function depending on the outputs of an ODE solver
@@
and the parameter gradient can be calculated by an integral Chen:2018
@@
Then we identify the shear response $\left(\chi_{\mathrm{re}},\chi_{\mathrm{im}}\right)$ with the state $x$ and the radial coordinate $z$ with the time $t$. In general, the function of metric $f(z)$ or the function of mass square $m^{2}(z)$ can be represented by a neural network with the parameter $\theta$.
Obviously, given the parameter $\theta$ is equivalent to giving the trial function of metric or mass square which specifies the Neural ODE. By invoking an ODE solver, one can integrate from IR to UV and predict the real part of shear viscosity. Using the adjoint sensitivity method, one can implement back-propagation and perform training. We depict the architecture of the current algorithm in figure 1. Note that it differs from a typical Neural ODE in two aspects: the data is not a time series but only relates to the output at the upper limit of the integration, and one component of the adjoint state vanishes.
@@ 3.2 Loss function and inductive bias
Our loss function is chosen as the log-cosh form LC2018 ; LC2020 ; LC2021 ; LC2022
@@
where $N$ denotes the size of data, $\bar{\eta}_{\mathrm{re}}$ represents the input data, and $\eta_{\mathrm{re}}$ is what the Neural ODE predicts. Note that the loss $\log\left[\cosh\left(x\right)\right]$ behaves like L1 loss $\left|x\right|$ far from the origin and L2 loss $x^{2}$ close to the origin. It is twice differentiable everywhere, making it suitable for optimizers like BFGS that require the Hessian Nocedal .
@@ 4 Holographic models
We will study three well-known holographic models without translation symmetry, all of which have homogeneous and isotropic backgrounds.
@@ 4.1 Massive gravity
The research of massive gravity (MG) has a long and winding history Pauli1939 ; Veltman1970 ; Zakharov1970 ; Vainshtein1972 ; Boulware1972 . The main theoretical interest is to explore whether there exists a self-consistent theory that has a massive graviton with spin 2. In cosmology, MG is considered as a candidate theory that explains the accelerated expansion of the universe by modifying Einstein gravity. In holography, massive gravity is the first analytically tractable model without translation symmetry Vegh1301
We will consider the dRGT massive gravity that is claimed to be ghost free deRham2010 ; deRham2011 ; deRham1401 . Its simplest version used in holography is described by the bulk action
@@
where $\mathcal{X}{{}^{\mu}}_{\nu}=\sqrt{g^{\mu\lambda}f_{\lambda\nu}}$ and $f_{\mu\nu}=\mathrm{diag}(0,0,1,1)$ is the reference metric. Note that we have set the AdS radius to one for convenience.
@@
Now we can generate the data for training. Inputting eq. (25) and eq. (28) into eq. (11), using the regular condition at the horizon, and setting the parameter $\alpha=1$, one can solve the flow equation and obtain the real part of shear viscosity on the boundary. We sample the frequency uniformly between 0.01 and 6 to generate 600 data points $\left[\omega,\eta_{\mathrm{re}}(\omega)\right]$, which are plotted in the left panel of figure 2.
We conduct two numerical experiments: fixing the mass to learn the metric and fixing the metric to learn the mass. The primary training results are shown in the middle and right panels of figure 2. Additionally, figure 6 and table 1 in Appendix B detail the hyperparameter tuning, minimum loss, and mean relative error (MRE).
Figure 2: Training data and performance of Neural ODEs for the MG model. The left panel shows the real part of the frequency-dependent shear viscosity, obtained by solving the flow equation with the true metric and mass. The middle and right panels illustrate the performance of Neural ODEs in learning the metric and mass respectively, where the curves represent the true values and the dots indicate the predictions.